OpenStudy (anonymous):

what is 11.75 x 8 1/3?

OpenStudy (anonymous):

11.75=47/4 8 1/3 =25/3 (47/4)*(25/3)=1175/12 in decimal form this is 97.92 exactly this is 97 and 11/12

OpenStudy (anonymous):

well i put 97.9 on my paper because it said to round to the nearest tenth, so i guess i was right